OpenStudy (turingtest):

Using the digit 7 five time, place the basic arithmetic signs in between the digits to obtain a total of 540

OpenStudy (turingtest):

without using 77, there mus be an operation between each number!

OpenStudy (anonymous):

is it possible to get it mathematically? or just guess? :D

OpenStudy (anonymous):

it should be impossible

OpenStudy (anonymous):

wild guess

OpenStudy (turingtest):

I'm thinking it is impossible: adding or subtracting 7/7=1 is the only way I can see to get to a number that is not divisible by 7, which only leaves three 7's left over, and the most we can make that into is 7*7*7=343 taking 540-(7/7)=539 and 539/7=77 hence it seems jimmyrep's answer is the only one possible, but I don't think this is a formal proof that it is impossible. -jimmyrep's answer was 77*7+(7/7)

OpenStudy (anonymous):

there r a total of 256 combinations, and as far as a human mind can think, none of these 256 is the right solution. but again its intitution which is being used here

OpenStudy (anonymous):

77*7 + 7/7